OpenStudy (anonymous):

I wanna make sure I got it right!! Simplify the expression: 2(3n+4)-2(2n+5)

OpenStudy (anonymous):

is it: 2n+18?

OpenStudy (1018):

2(3n+4) = 6n + 8 -2(2n+5) = -4n+5 = 6n + 8 - 4n + 5

OpenStudy (1018):

now if you simplify the last equation it would be: 2n-2

OpenStudy (1018):

sorry wait it shoulde be -4n -10 for the 2nd equation above

OpenStudy (1018):

simplifying 2n-2 further it would be 2(n-1)

OpenStudy (anonymous):

thanks so much:)